A=1/2bh A=38sqcm b=8cm solve for h. please show all work
Answer by stanbon(75887)

A=1/2bh A=38sqcm b=8cm solve for h. please show all work
-------------------
Substitute and solve for "h":
38 = (1/2)*8*h
---
38 = 4h
h = 38/4
h = 19/2 = 9.5 cm
======================
Cheers,
Stan H.